OpenStudy (anonymous):

I have a huge problem I need help with.. "Match the number and type of solutions for the following two functions in standard form: f(x) = x^2 + 6x – 16 g(x) = x^2 +6x + 1 Match the following descriptions of the solutions to each of the functions above. Hint: they each have their own match. Two real irrationals solutions. Two real rational solutions. Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. Find the solutions of g(x). Show each step."

OpenStudy (anonymous):

For f(x)=x^2+6x-16, I got "x=-8,2" Am I doing it right?

OpenStudy (anonymous):

For g(x)=x^2+6x+1, I got "-3,-8"

OpenStudy (anonymous):

@Nateyboo